There are 18 more sophomores than juniors in an algebra class. If there are 78 students in this class, find the number of sophomores and the number of
juniors in the class.

Answer by macston(5194)

.
If all students are sophomores or juniors:
.
J=number of juniors; S=number of sophomores=J+18
.
J+S=78
J+J+18=78
2J=60
J=30
ANSWER 1: There are 30 juniors in the class.
S=J+18=30+18=48
ANSWER 2: There are 48 sophomores in the class.
.
CHECK:
J+S=78
30+48=78
78=78